OpenStudy (anonymous):

Log8 cube root 64 find the exact value = log8 4 = log10 4/log10 8?

geerky42 (geerky42):

This?\[\Large \log_8 \sqrt[3]{64}\]

OpenStudy (anonymous):

Yes

geerky42 (geerky42):

Oh ok, you can rewrite root to exponent, like this: \[\Large \log_8 \sqrt[3]{64} \Rightarrow \log_8 64^{\frac{1}{3}}\]

geerky42 (geerky42):

We need to do something to 1/3

OpenStudy (anonymous):

I can then move the 1/3 to the front

geerky42 (geerky42):

Right!

OpenStudy (anonymous):

Can. I then write 1/3 log8 4^3?

geerky42 (geerky42):

hmm, you can do it, or you could just write \(\Large \dfrac{1}{3}\log_8 8^2\)

geerky42 (geerky42):

it's easier to solve this way. Move 2 to front, like this: \[\Large \dfrac{2}{3} \log_8 8\]

OpenStudy (anonymous):

I see. , my answer is 2/3

geerky42 (geerky42):

right.

OpenStudy (anonymous):

Thank you so much!

geerky42 (geerky42):

glad I helped!